# Robust Market Interventions

### v1

@@ 5.1. Generating the true and observed market state
5.2. Interventions
We focus on the following intervention rule: Recover the eigenvector associated to the largest eigenvalue, in absolute terms, of the estimated Slutsky matrix $\widehat{\bm{D}}$. Intervene to subsidize firms in proportion to this eigenvector. This intervention aligns with the intervention rule behind our Theorem 1.202020In this example, for low values of $\gamma$ the largest eigenvalue of $\bm{D}$ is sufficiently well-separated from all other eigenvalues and, consequently, the authority can use the first eigenvector of $\widehat{\bm{D}}$ as a good approximation of $\bm{u}^{1}$ (see the illustrative discussion in Section 4.2.2).
@@ 5.3. Evaluation of interventions
Figure 4. The effects of first eigenvector intervention on changes in consumer surplus (panel a) and changes in producer surplus (panel b) as a function of $\gamma$ in a market described in Section 5.1. For each value of $\gamma$, we plot the median (in blue), 5th and 95th percentiles associated with 3000 market state realizations.
Figure 4 shows a sharp transition in the performance of the intervention. For $\gamma$ less than roughly $0.7$ the true market state has very large eigenvalues and so the authority can use the estimate of $\bm{D}$ to precisely identify the underlying main eigenvector(see Figure 1 for $\gamma=0.3$.) Note also that products in category $3$ are the ones who are most substitutable with other products in categories $1$ and $2$ and so they are highly represented in the first eigenvector (i.e., they have a high eigenvector centrality). This implies that the estimated first eigenvector is sufficiently correlated to the status quo market quantity. Hence, for low $\gamma$, the true market state has recoverable structure. This allows the authority to implement interventions that robustly have the pass-through properties characteristic of high-eigenvalue eigenvectors—negligible impact on prices and hence on consumers, along with an effect on producer surplus equal to twice the authority’s spending , which is normalized to one unit.
@@
When the authority can estimate the first eigenvector accurately, the first eigenvector intervention turns out to subsidize products in block $3$ and tax all the other products. Subsidizing products in block 3 leads to a decrease in their price and hence an increase in their demand. Taxing products in blocks 1 and 2 leads to an increase of the price of these products, and hence an increase in the demand of products in block 3. Combining the effects, the intervention leads to a relatively high increase in the demand of good 3, while keeping prices roughly constant. As a result, both producer and total surplus increase dramatically without sizeable changes in consumer surplus.
@@ 6. Tightness of the main result
We have established that if demand has recoverable structure, the authority can robustly achieve the maximum possible total surplus per dollar spent subject to the constraint that consumers are not harmed. The associated intervention rule boosts production with minimal price changes, resulting in all efficiency gains being captured by firms.
@@
In each case, 2 describes the limits on what can be achieved by an authority with noisy information. It also notes that these limits really are about information: part (ii) of each case states that an omniscient authority would not be subject to the same limitation.
In more detail, Part 1 of 2 tells us that we may not be able to design interventions that robustly increase total surplus $\dot{W}=\dot{C}+\dot{P}-\dot{S}$. (Since, by 1 we know this can be done under the recoverable structure assumption, our construction must lack recoverable structure.) Intuitively, in this case the information about the market state learned from the signal can be very imprecise and, therefore, there are market states in which any intervention will lead to undesirable outcomes. The proof constructs a set of market states such that, with an uninformative signal, for any intervention there is a market state with $\dot{W}<0$. The basic idea is to use the total surplus decomposition:
@@
Part 2 of 2 tells us that even if the market state has recoverable structure, it may be impossible to design interventions that robustly increase total surplus and allow consumers to capture some of the resulting efficiency gains. 2 tells us that to achieve such an outcome, the intervention must project onto some $\bm{u}^{\ell}$ where $\lambda_{\ell}$ is not too large, since only those eigenvectors have nonvanishing pass-through to consumer surplus. However, the noisy observation of $\widehat{\bm{D}}$ and $\bm{q}^{0}$ gives very noisy estimates of the constituents of (10) corresponding to these eigenvectors. So, by targeting them, there is a sizeable chance (at least in some environments) that the policy will have negative consequences for the consumers.
@@ 7. Discussion and concluding remarks
We have developed a theory of robust interventions in large oligopolies. We identify a condition on demand under which an authority can robustly increase the total surplus per dollar spent as much as would be possible under perfect information subject to the constraint that consumers are not harmed. The methodological contribution is the development of spectral methods to analyze pass-through in oligopolies, and the application of these methods to gain leverage on statistical problems about oligopolies observed with noise.
@@ 7.1. Marketplace big data in practice
In our model, the authority has “big data” about the Slutsky matrix $\bm{D}$ that may not allow precise estimation of any pairwise demand interactions or hedonic model parameters. This seems natural for markets with a large and changing collection of goods, such as those hosted on large online platforms. Such platforms collect immense amounts of data—about consumer browsing behavior, timing of purchases, consideration sets, etc.—and apply machine learning techniques to these data to form informative but imperfect estimates of interactions among various goods (Athey, 2018; Wager and Xu, 2021; Cai and Daskalakis, 2022; Bajari, Burdick, Imbens, Masoero, McQueen, Richardson, and Rosen, 2023). This is modeled by our notion of a market signal. [^22] The distributional properties of the matrix $\bm{E}$ capturing the errors in these estimates would depend on the application, as would the conditions for $\bm{E}$ to have small enough norm for recoverable structure to be present. It would be interesting to investigate these issues in specific applications. Our paper offers an approach for calculating suitable statistics that suffice for effective interventions despite the noise in this signal.
@@
Our approach contrasts with a standard one in empirical industrial organization, where markets are defined tightly so that each contains only a relatively small number of similar goods with strong demand interactions, and then a small number of hedonic parameters and demand elasticities are precisely estimated. That approach would correspond in our notation to a very precise signal (i.e., an error matrix $\bm{E}$ with small norm) and a small number of goods.
@@
In terms of interventions, online marketplaces are an appealing setting for our results because the authorities in charge of them can finely target policies that function as taxes and subsidies, including, e.g., commission rates, discount coupons, free advertising, etc.—and regularly experiment with such perturbations. It is worth noting, however, that the policies our analysis recommends need not be specific to individual products. This is because when we take a large matrix (in our case, the Slutsky matrix) that reflects relationships among units (in our case, products) and look at the eigenvectors with large eigenvalues, the coordinates of those eigenvectors typically yield low-dimensional embeddings capturing substantively natural categories (Chen, Chi, Fan, and Ma, 2021). For instance, in our Example 1, the top two eigenvectors are sufficient to recover the blocks to which the goods belong. Relatedly, spectral clustering analyses based on the top few eigenvectors sort items into natural “similarity” classes, where similarity is defined by having similar relationships to other classes (Spielman and Teng, 1996). Once again, the spectral statistics used in these techniques tend to pick up interpretable “broad” features of the products, rather than idiosyncracies specific to individual products. As a result, the policies our interventions recommend—which project all variation onto these eigenvectors—will often be close to a policy that depends mostly on category—e.g., a subsidy on the sale of smartphones along with a tax on certain types of accessories. Though the policies will not be perfectly regular (note the irregularities of Panel D of Figure 1) the above observations lead us to conjecture that an authority constrained to design policies that discriminate only at a coarse product level could, under natural assumptions, achieve a substantial amount of the gains of our policies. We leave these interesting considerations to future work.
@@ 7.2. Relationship with hedonic utility models
Recent work by Pellegrino (2021) and Ederer and Pellegrino (2021) uses an oligopoly model to empirically quantify the evolution of market power. The relationship of our model to their work sheds light on the types of empirical models that can capture recoverable structure. In Pellegrino (2021), the model of demand is hedonic, in the spirit of Lancaster (1966): the household’s utility is additively separable in the contributions of various characteristics, and a product provides a bundle of these characteristics. The Slutsky matrix $\bm{D}$ derived from this demand model can be expressed as a transformation of the cosine similarity matrix of products’ characteristics, which was estimated for a large set of consumer goods by Hoberg and Phillips (2016) using text data. [^23] Pellegrino (2021) and Ederer and Pellegrino (2021) consider quantity competition, but the Slutsky matrix does not depend on this choice. We have calculated that in the Slutsky matrix derived this way, the eigenvalues are all small and the recoverable structure condition fails. [^24] There are more than $n>3000$ products in the data and for the case of i.i.d. noise we would require that the largest eigenvalue is considerably larger than $b(n)=\sqrt{n}\approx 54$; this fails as largest eigenvalue of $\bm{D}$ in absolute value is about $2$. It is useful to reflect on why this is the case.
@@
In the model of Pellegrino (2021), a simple calculation shows that it is not possible for the Slutsky matrix to have large eigenvalues. [^25] The (un-normalized) Slutsky matrix in Pellegrino (2021) is $-\bm{B}^{-1}$, where $\bm{B}=\bm{I}+\alpha(\bm{\Sigma}-\bm{I})$. The matrix $\bm{\Sigma}$ is positive semidefinite because it can be written as $\bm{V}^{\mathsf{T}}\bm{V}$, where the columns of $\bm{V}$ are the characteristic vectors of various products. Thus all eigenvalues of $\bm{B}$ are real numbers bounded below by $1-\alpha$, and all eigenvalues of $-\bm{B}^{-1}$ are at most $1/(1-\alpha)$ in magnitude. Pellegrino uses the value $\alpha=0.12$, which prevents any eigenvalue from exceeding $1.13$. We do not work with exactly the same Slutsky matrix because of the normalization in Section A.1. Its eigenvalues are a bit different, but they can still be bounded by a constant. Numerically we see that the normalization makes little difference. For an economic intuition, note that in the Lancaster (1966) type of model, the “direct” relationship between any pair of goods is substitution. With substitution, if some demand is diverted from one good due an increase in its price, the sum of effects on all substitute goods is bounded, since, loosely speaking, the demand gained by these other goods must come out of the demand lost by the more expensive one. So the sum of positive entries in $\bm{D}$ corresponding to this effect is bounded. This in turn places bounds on any complementarities in the Slutsky matrix. [^26] Note that complementarity—demand of one good is decreasing in the price of the other—can arise in Pellegrino (2021) model. This happens through indirect effects: the substitute of my substitute can be my complement. However, since the “direct” substitution effect is bounded in magnitude, so are the indirect consequences. In short, in a hedonic model where the basic force is substitution, overall spillovers are bounded, and the fact that $\bm{D}$ has no large eigenvalues is the mathematical manifestation of this.
@@
Quite different behavior is seen in models where utility arises directly out of the consumption of complementary goods, and such complementarities play a central role in our examples of recoverable structure. A leading practical example comes from the use of computers: a consumer’s utility from a computer depends on the hardware, the operating system, and the applications. Two firms selling distinct components—a hardware device and an operating system, for instance—are supplying complementary goods; on the other hand, two firms selling the same component (say, operating systems) are supplying substitute goods (Matutes and Regibeau, 1988, 1992). The illustrative example 1 we develop in Section 3 shows how Slutsky matrices with large eigenvalues arise naturally in such settings. [^27] The complementarities there happen to be within-category, but that is not important for our point here. But, as we have seen, it is impossible to produce the same patterns in models of the Lancaster (1966) type, because they cannot have large eigenvalues; one would need to add terms corresponding to the idea that some characteristics provide more value when enjoyed together. There is a simple economic intuition for why such complementarities are more likely to produce recoverable structure: when the price of one good goes down, then all its complements can become more demanded by comparable nonvanishing amounts if they are not substitutes for one another. This leaves the kinds of clusters of nonvanishing entries in $\bm{D}$ which are the hallmark of recoverable structure.
@@
In summary, direct complementarities of this sort seem practically important, as a natural source of the recoverable structure that plays a central role in our results. We hope these observations will motivate further empirical research on the structure of large-scale oligopoly models with complementary goods.
@@ 7.3. Games on networks
One can view our exercise as a special case of an intervention, under noisy information, in a game among a large number of agents. In our case, the game is a standard oligopoly pricing model. Note that, under the assumption of linear demand, the pricing game can be seen as a network game with linear best replies where the Slutsky matrix is the network. Our analysis shows that if the oligopoly exhibits a recoverable structure, then there are robust interventions that maximise changes in producer surplus. The methods we have developed can be extended to other settings. For example, in a public goods setting, interventions aim to realign private marginal returns with social marginal returns. The literature has developed tools to understand how to do this when the authority has precise information on the spillovers causing the underprovision of public goods. [^28] See, for instance, Bramoullé et al. (2014) and Elliott and Golub (2019). However, we know little about the design of interventions under noisy information about such externalities. Similarly, in contracting for teams under moral hazard, network methods have recently been developed for locally perturbing contracts to achieve better outcomes for a principal (Dasaratha et al., 2024). But it is a considerable challenge to extend these results to the realistic case where the strategic interactions among members of an organization are only imperfectly known. General games will lack some of the structure we have taken advantage of. This includes the properties of the spillovers structure coming from a symmetric, positive semidefinite Slutsky matrix. So there are challenges to overcome in extending our results. We hope this paper stimulates research in these exciting new directions.

### v2

@@ 5.2. Evaluation of interventions
Figure 4. The effects of a first eigenvector intervention on changes in consumer surplus (panel a) and changes in producer surplus (panel b) as a function of $\gamma$ in a market described in Example 2. For each value of $\gamma$, we plot the median (in blue), 5th, and 95th percentiles associated with 3000 market state realizations.
Figure 4 shows a sharp transition in the performance of the intervention. For $\gamma$ less than roughly $0.7$ the true market state has very large eigenvalues and so the authority can use the estimate of $\bm{D}$ to precisely identify the underlying main eigenvector(see Figure 2 for $\gamma=0.3$.) Note also that products in category $3$ are the ones that are most substitutable with other products in categories $1$ and $2$ and so they are highly represented in the first eigenvector (i.e., they have a high eigenvector centrality). This implies that the estimated first eigenvector is sufficiently correlated with the status quo market quantity. Hence, for low $\gamma$, the true market state has recoverable structure. This allows the authority to implement interventions that robustly have the pass-through properties characteristic of high-eigenvalue eigenvectors—negligible impact on prices and hence on consumers, along with an effect on producer surplus equal to twice the authority’s spending , which is normalized to one unit.
@@
When the authority can estimate the first eigenvector accurately, the first eigenvector intervention turns out to subsidize products in block $3$ and tax all the other products. Subsidizing products in block 3 leads to a decrease in their price and hence an increase in their demand. Taxing products in blocks 1 and 2 leads to an increase of the price of these products, and hence an increase in the demand of products in block 3. Combining the effects, the intervention leads to a relatively high increase in the demand of good 3, while keeping prices roughly constant. As a result, both producer and total surplus increase dramatically without sizable changes in consumer surplus.
@@ 6. Tightness of the main result
We have established that if demand has recoverable structure, the authority can robustly achieve the maximum possible total surplus per dollar spent subject to the constraint that consumers are not harmed. The associated intervention rule boosts production with minimal price changes, resulting in firms capturing all efficiency gains.
@@
In each case, 3 describes the limits on what can be achieved by an authority with noisy information. It also notes that these limits really are about information: part (ii) of each case states that an omniscient authority would not be subject to the same limitation.
In more detail, Part 1 of 3 tells us that we may not be able to design interventions that robustly increase total surplus $\dot{W}=\dot{C}+\dot{P}-\dot{S}$. (Since, by 1 we know this can be done under the recoverable structure assumption, our construction must lack recoverable structure.) Intuitively, in this case the information about the market state learned from the signal can be very imprecise and, therefore, there are market states in which any intervention will lead to undesirable outcomes. The proof constructs a set of market states such that, with an uninformative signal, for any intervention there is a market state with $\dot{W}<0$. The basic idea is to use the total surplus decomposition:
@@
Part 2 of 3 tells us that even if the market state has recoverable structure, it may be impossible to design interventions that robustly increase total surplus and allow consumers to capture some of the resulting efficiency gains. 3 tells us that to achieve such an outcome, the intervention must project onto some $\bm{u}^{\ell}$ where $\lambda_{\ell}$ is not too large, since only those eigenvectors have nonvanishing pass-through to consumer surplus. However, the noisy observation of $\widehat{\bm{D}}$ and $\bm{q}^{0}$ gives very noisy estimates of the constituents of (13) corresponding to these eigenvectors. So, by targeting them, there is a substantial chance (at least in some environments) that the policy will have negative consequences for consumers.
@@ 7. Discussion and concluding remarks
We have developed a theory of robust interventions in large oligopolies. We identify a condition on demand under which an authority can robustly increase the total surplus per dollar spent as much as would be possible under perfect information subject to the constraint that consumers are not harmed. The methodological contribution lies in developing spectral methods to analyze pass-through in oligopolies, and applying these methods to gain leverage on statistical problems about oligopolies observed with noise.
@@ 7.1. Marketplace big data in practice
In our model, the authority has “big data” about the Slutsky matrix $\bm{D}$ that may not allow precise estimation of any pairwise demand interactions or hedonic model parameters. This is natural for markets with a large and changing collection of goods, such as those hosted on large online marketplaces. Such marketplaces collect immense amounts of data—about consumer browsing behavior, timing of purchases, consideration sets, etc.—and apply machine learning techniques to these data to form informative but imperfect estimates of interactions among various goods (Athey, 2018; Wager and Xu, 2021; Cai and Daskalakis, 2022; Bajari, Burdick, Imbens, Masoero, McQueen, Richardson, and Rosen, 2023). This is modeled by our notion of a market signal. [^28] The distributional properties of $\bm{E}$ describing the errors in these estimates would depend on the application, as would the conditions for $\bm{E}$ that would bound its norm. It would be interesting to investigate these issues in specific applications. Our paper offers an approach for calculating suitable statistics that suffice for effective interventions despite the noise in this signal.
@@
Our approach contrasts with a standard one in empirical industrial organization, where markets tend to be defined tightly so that each contains only a relatively small number of similar goods with strong demand interactions, and then a small number of hedonic parameters and demand elasticities are precisely estimated. That approach would correspond in our notation to a very precise signal (i.e., an error matrix $\bm{E}$ with small norm) and a small number of products.
We have focused on total market surplus as a canonical objective. But we also saw that our robust intervention maximizes the change in total surplus by increasing producer surplus while holding consumer surplus constant. In this sense, the intervention maximizes the increase in producer surplus under the constraint that consumers do not lose—a reasonable objective for an operator of a marketplace that collects revenue proportional to sellers’ profits. We leave the study to other objective functions for the authority to future research.
@@
Our model permits flexible marginal interventions. This modeling choice is suited to online marketplaces, because the operators in charge of them can finely target policies that function as taxes and subsidies, including commission rates, discount coupons, free advertising, etc.—and regularly experiment with such perturbations. It is worth noting, however, that the policies our analysis recommends need not be specific to individual products. This is because when we take a large matrix (in our case, the Slutsky matrix) reflecting relationships among units (in our case, products) and look at eigenvectors with large eigenvalues, the coordinates of those eigenvectors typically yield low-dimensional embeddings capturing substantively natural categories (Chen, Chi, Fan, and Ma, 2021). For instance, in our Example 1, the top two eigenvectors are sufficient to recover the blocks to which the goods belong. Relatedly, spectral clustering analyses based on the top few eigenvectors sort items into natural “similarity” classes, where similarity is defined by having similar relationships to other classes (Spielman and Teng, 1996). Once again, the spectral statistics used in these techniques tend to pick up interpretable “broad” features of the products, rather than idiosyncrasies specific to individual products. As a result, the policies our interventions recommend—which project all variation onto these eigenvectors—will often be close to a policy that depends mostly on category—e.g., a subsidy on smartphones along with a tax on certain types of accessories. Though the policies will not be perfectly regular (note the irregularities of Panel D of Figure 2) the above observations lead us to conjecture that an authority constrained to design policies that discriminate only at a coarse product level could, under natural assumptions, achieve a substantial amount of the gains of our policies. We leave these interesting considerations to future work.
Lastly, we mention the problem of predicting the effects of general perturbations to markets in settings such as ours. This is related to, but distinct from, the problem we have studied. The problem of robust intervention is importantly easier, because the authority chooses the perturbation to make a prediction about. Nevertheless, the spectral decomposition of intervention effects appears likely to be useful for descriptive comparative statics in cases where a cost shock $\bm{\sigma}$ or other change is exogenous.
@@ 7.2. Relationship with hedonic utility models
Recent work by Pellegrino (2021) and Ederer and Pellegrino (2021) uses an oligopoly model to empirically quantify the evolution of market power. The relationship between our model and their work sheds light on the types of empirical models that can capture recoverable structure.
@@
In Pellegrino (2021), the model of demand is hedonic, in the spirit of Lancaster (1966): the household’s utility is additively separable in the contributions of various characteristics, and a product provides a bundle of these characteristics. The Slutsky matrix $\bm{D}$ derived from this demand model can be expressed as a transformation of the cosine similarity matrix of products’ characteristics, which Hoberg and Phillips (2016) estimated for a large set of consumer goods using text data. [^29] Pellegrino (2021) and Ederer and Pellegrino (2021) consider quantity competition, but the Slutsky matrix does not depend on this choice. We have calculated that in the Slutsky matrix derived this way, the eigenvalues are all small and the recoverable structure condition fails. [^30] There are more than $3000$ products in the data and for the case of i.i.d. noise we would require that the largest eigenvalue is considerably larger than $b(n)=\sqrt{n}\approx 54$; this fails as largest eigenvalue of $\bm{D}$ in absolute value is about $2$. It is useful to reflect on why this is the case.
@@
In the model of Pellegrino (2021), a simple calculation shows that it is impossible for the Slutsky matrix to have large eigenvalues. [^31] The (un-normalized) Slutsky matrix in Pellegrino (2021) is $-\bm{B}^{-1}$, where $\bm{B}=\bm{I}+\alpha(\bm{\Sigma}-\bm{I})$. The matrix $\bm{\Sigma}$ is positive semidefinite because it can be written as $\bm{V}^{\mathsf{T}}\bm{V}$, where the columns of $\bm{V}$ are the characteristic vectors of various products. Thus all eigenvalues of $\bm{B}$ are real numbers bounded below by $1-\alpha$, and all eigenvalues of $-\bm{B}^{-1}$ are at most $1/(1-\alpha)$ in magnitude. Pellegrino uses the value $\alpha=0.12$, which prevents any eigenvalue from exceeding $1.13$. We do not work with exactly the same Slutsky matrix because of the normalization in Section A.1. Its eigenvalues are a bit different, but they can still be bounded by a constant by elaborating this argument. Numerically we see that the normalization makes little difference. For an economic intuition, note that in the Lancaster (1966) type of model, the “direct” relationship between any pair of goods is substitution. With substitution, if some demand is diverted from one good due to an increase in its price, the total effect on all substitute goods is bounded, since, loosely speaking, the demand gained by these other goods must come out of the demand lost by the more expensive one. This bounds the sum of positive entries in $\bm{D}$ corresponding to this effect, which in turn bounds any complementarities in the Slutsky matrix. [^32] Note that complementarity (where one good’s demand decreases in the price of the other) can arise in Pellegrino (2021) model. This happens through indirect effects: the substitute of my substitute can be my complement. However, since the “direct” substitution effect is bounded in magnitude, so are the indirect consequences. In essence, in a hedonic model where the basic force is substitution, overall spillovers remain bounded, and the fact that $\bm{D}$ has no large eigenvalues is the mathematical manifestation of this.
@@
Quite different behavior emerges in models where utility arises directly from consuming goods together, and such complementarities are central to our examples of recoverable structure. A leading practical example comes from the use of computers: a consumer’s utility from a computer depends on the hardware, operating system, and applications. Two firms selling distinct components—a hardware device and an operating system, for instance—supply complementary goods, while two firms selling the same component (say, operating systems) supply substitute goods (Matutes and Regibeau, 1988, 1992). Our illustrative Example 2 in Section 5 shows how Slutsky matrices with large eigenvalues arise naturally in such settings. [^33] The complementarities there happen to be within-category, but that is not important for our point here. But, as we have seen, it is impossible to produce the same patterns in models of the Lancaster (1966) type, because they cannot generate large eigenvalues; one would need to incorporate terms reflecting that some characteristics provide greater value when enjoyed together. There is a straightforward economic intuition for why such complementarities more readily produce recoverable structure: when one good’s price decreases, all its complements can experience comparable nonvanishing increases in demand. This creates the clusters of nonvanishing entries in $\bm{D}$ that are the hallmark of recoverable structure.
@@
In summary, direct complementarities seem practically important and can naturally yield the recoverable structure central to our results. We hope these observations will motivate further empirical research on the structure of large-scale oligopoly models with complementary goods.
@@ 7.3. Games on networks
One can view our exercise as a special case of an intervention, under noisy information, in a game among a large number of agents. In our case, the game comes from a standard oligopoly pricing model. Under the assumption of linear demand, the pricing game can be seen as a network game with linear best replies where the Slutsky matrix defines the network. Our analysis shows that if the oligopoly exhibits recoverable structure, then there are robust interventions for particular economic objectives. The methods we have developed can be extended to other settings. For example, in a public goods setting, interventions would aim to realign private marginal returns with social marginal returns. The literature has developed tools to understand how to do this when the authority has precise information on the spillovers causing the underprovision of public goods. [^34] See, for instance, Bramoullé et al. (2014). However, we know little about designing interventions under noisy information about such externalities. Similarly, in contracting for teams under moral hazard, network methods have recently been developed for locally perturbing contracts to achieve better outcomes for a principal (Dasaratha et al., 2024). But it is a considerable challenge to extend these results to the realistic case where the strategic interactions among members of an organization are only imperfectly known. General games will lack some of the structure we have leveraged, including the properties of the spillovers structure coming from a symmetric, positive semidefinite Slutsky matrix. So there are challenges to overcome in extending our results. We hope this paper stimulates research in these directions.